Please help me with this problem:
Given the function f defined by f(x)=cos(x)-cos^2(x) (for -pi being less than or equal to x which is less than or equal to pi).
Find the x-intercepts of the graph of f.

Answer by stanbon(75887)

Given the function f defined by f(x)=cos(x)-cos^2(x) (-pi <= x <= pi).
Find the x-intercepts of the graph of f.
---------------------
Let f(x)=0
cos^2(x)-cos(x) = 0
[cos(x)][cos(x)-1]=0
cosx = 0 or cosx = 1
x = pi/2 or x = 0
==============
Cheers,
Stan H.